Jim is 3 times as old as Jane was 2 years ago. in 10 years Jane will be as old as Jim is now. how old are they now?
Answer by mananth(16946)

Jane's age now = x
2 years ago his age = x-2
Jim is now 3(x-2)
Present ages Jane =x
Jim = 3(x-2)
10 years later
Jane's age = x+10
x+10=3(x-2)
x+10=3x-6
3x-x= 10+6
2x= 16
/2
x= 8
Jane's age now = 8 years
Jims age = 3(x-2)
=3(8-2)
=18
Jim's age = 18 years now